The admission fee at small fair is $1.50 for children and $4.00 for adults. On a certain day, 2,200 people enter the fair and $5050 is collected. How many children and how many adults attended?
Answer by nerdybill(7384)

The admission fee at small fair is $1.50 for children and $4.00 for adults. On a certain day, 2,200 people enter the fair and $5050 is collected. How many children and how many adults attended?
.
Let A = number of adults
then because a total of 2200 people entered:
2200-A = number of children
.
Since the total collected was 5050:
5050 = 4A + 1.50(2200-A)
5050 = 4A + 3300 - 1.50A
5050 = 2.5A + 3300
1750 = 2.5A
700 = A (number of adults)
.
number of children:
2200-A = 2200-700 = 1500